You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update: Weekend update 14 November 2014

The cyclotron is running with BL2A and BL2C4 both online. It’s not clear how long it will take for the BL1A vacuum to recover to where it can also be brought online; the current estimate is some time tomorrow (Saturday).

Beam delivery update: Maintenance underway 13 November 2014

The cyclotron is off for maintenance and will remain off overnight. It’s not clear at this point when either the cyclotron or BL1A will be ready for operation.

Beam delivery update: Cyclotron defrost tomorrow 12 November 2014

The cyclotron’s currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~40 µA to BL2C4, while at ISAC 8Li is still being delivered to βNMR/NQR. There are three shifts of maintenance scheduled starting tomorrow to accommodate a defrost of the cyclotron cryopanels and cryogenics system.

Beam delivery update: Lab closed tomorrow; long maintenance Thursday 10 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~50 µA to BL2C4. 8Li is being delivered to βNMR/NQR at ISAC for materials science experiments. This is expected to continue with only minor changes until Thursday morning.

Beam delivery update: Weekend edition 07 November 2014

The cyclotron is delivering ~110 µA to BL1A and 50 µA to BL2A. We’re scheduled to run with high currents in all three beamlines over the weekend and into next week.

Beam delivery update 06 November 2014

The BL2A current was reduced to 50 µA earlier today and BL2C4 is currently running at 80 µA, but BL1A is still running at ~110 µA and 8Li is still being delivered to βNMR/NQR at ISAC.

Beam delivery update 05 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue through the weekend.

Beam delivery update: ISAC maintenance tomorrow 04 November 2014

The cyclotron’s currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue until ~08:30 tomorrow morning at which time BL2A will be defined off for a target module move.

Beam delivery update: No maintenance tomorrow; ISAC maintenance Wednesday 03 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4 (limited by the resistivity of the STF cooling water). This is expected to continue until Wednesday morning.

Beam delivery update 31 October 2014

The cyclotron is still delivering ~115 µA of beam to BL1A, 60 µA to BL2A, and ~95 µA to BL2C4. The BL2A current might still be adjusted, but otherwise we expect to run like this through the weekend.

Pages

Subscribe to Group News